Two standard six-sided dice are rolled. The probability that the sum of the two dice is 9 or larger is
A.1/3 B.1/4 C. 1/6 D. 1/9 E. 5/18

Answer by stanbon(57387)

Two standard six-sided dice are rolled. The probability that the sum of the two dice is 9 or larger is
---
List the "successes":
3/6;4/5;5/4;6/3
4/6;5/5;6/4
5/6;6/5
6/6
Count them: 10
P(sum >=9) = 10/36 = 5/18
-----------
Cheers,
Stan H.


A.1/3 B.1/4 C. 1/6 D. 1/9 E. 5/18